Equation of a Tangent Line: Find the standard (slope-intercept form) equation of the tangent line to the following functions at the specified points. (Show your complete solution, use Method 1)

  1. ๐‘“(๐‘ฅ) = ๐‘ฅ2 โˆ’ 4๐‘ฅ+1 at point (0,1)
  2. ๐‘“(๐‘ฅ) = ๐‘ฅ3+ 2 at point (1,3)
  3. ๐‘“(๐‘ฅ) = 2๐‘ฅ4+ 3๐‘ฅ3 โˆ’ 2๐‘ฅ + 7 at point (-1,8)
  4. ๐‘“(๐‘ฅ) = โˆš2๐‘ฅ + 7 ๐‘Ž๐‘ก ๐‘ฅ =1
1
Expert's answer
2022-05-17T11:15:43-0400

1.


"f'(x)=(x^2-4x+1)'=2x-4"

"slope=m=f'(0)=2(0)-4=-4"

"y-1=-4(x-0)"

The standard (slope-intercept form) equation of the tangent line is


"y=-4x+1"

2.


"f'(x)=(x^3+2)'=3x^2"

"slope=m=f'(1)=3(1)^2=3"

"y-3=3(x-1)"

The standard (slope-intercept form) equation of the tangent line is


"y=3x"



3.


"f'(x)=(2x^4+ 3x^3 \u2212 2\ud835\udc65 + 7)'=8x^3+9x^2-2"

"slope=m=f'(-1)=8(-1)^3+9(-1)^2-2=-1"

"y-8=-(x-(-1))"

The standard (slope-intercept form) equation of the tangent line is


"y=-x+7"

4.


"f'(x)=(\sqrt{2x+7})'=\dfrac{1}{\sqrt{2x+7}}"

"slope=m=f'(1)=\dfrac{1}{\sqrt{2(1)+7}}=\dfrac{1}{3}"

"f(1)=\sqrt{2(1)+7}=3"

"y-3=\dfrac{1}{3}(x-1)"

The standard (slope-intercept form) equation of the tangent line is


"y=\dfrac{1}{3}x+\dfrac{8}{3}"